# Interpretable Pre-Trained Transformers for Heart Time-Series Data

### v1

@@ 8 Conclusion
This work has conclusively demonstrated that GPT-like models, trained to predict the next token in periodic physiological time series, are fully interpretable in their operation. This has been illustrated through aggregate attention maps which are logical for the task of predicting the next token and through the clustering of different PPG and ECG features in vector space. This is further shown through individual attention heads that respond strongly to specific features such as the dicrotic notch in PPG or the P-wave in ECG. Moreover, we have shown that this interpretability is carried forward when fine-tuning for the classification of atrial fibrillation, where attention shifts to regions in the input context that most indicate the presence of the arrhythmia. This work represents a step forward in the explainability of large transformer networks when applied to healthcare settings.

### v2

@@ 6 Conclusion
This work has conclusively demonstrated that GPT-like models, trained to predict the next token in periodic physiological time series, are fully interpretable in their operation. This has been illustrated through aggregate attention maps which are natural for the task of predicting the next token and through the clustering of different PPG and ECG features in vector space. This has been further shown through individual attention heads that correspond strongly to specific features such as the dicrotic notch in PPG or the P-wave in ECG. Moreover, we have shown that this interpretability is carried forward when fine-tuning for the classification of atrial fibrillation, where attention shifts to regions in the input context that most indicate the presence of the arrhythmia. This work represents a step forward in the explainability of large transformer networks when applied to healthcare settings.